OpenStudy (anonymous):

solve by the elimination method: 7x + 5y = -18 -7x + y = 30 What is the solution of the system?

OpenStudy (anonymous):

again, always try to eliminate a variable first. as u can see, it's faster if u eliminate y out from the second eqn.

OpenStudy (vishweshshrimali5):

I can help you again

OpenStudy (vishweshshrimali5):

See add both the euations

OpenStudy (anonymous):

Ok. thanks

OpenStudy (vishweshshrimali5):

you will get

OpenStudy (vishweshshrimali5):

6y = -18 + 30 => 6y = 12 => y = 2 Thus substitute this value in first equation

OpenStudy (vishweshshrimali5):

7x + 5y = -18 => 7x + 10 = -18 => 7x = -28 => x = -4 Thus , x = -4 and y = 2